OpenStudy (anonymous):

Use the quadratic formula to solve x2 + 7x + 8 = 0. Estimate irrational solutions to the nearest tenth. A. {–12, 5} B. {–11.1, –2.9} C. {–5.6, –1.4} D. {–3.9, 12.1}

OpenStudy (welshfella):

x = -7 +/- sqrt(7^2 - 4*1*8) --------------------- 2

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

are you confident in working that out?

OpenStudy (anonymous):

i barley started to learn this, so i you can walk through it with me?

OpenStudy (welshfella):

ok well use ur calculator its probably easier to work out sqrt(7^2 - 4*1*8) first this = sqrt(49 - 32) = sqrt 17

OpenStudy (welshfella):

can you find sqrt17 on ur calculator?

OpenStudy (anonymous):



OpenStudy (welshfella):

yes

OpenStudy (anonymous):

4 is the closest

OpenStudy (welshfella):

you need to be more accurate than that - find it to the nearest hundredth

OpenStudy (anonymous):

my calcuator doesnt do that!

OpenStudy (welshfella):

oh ok - are you on a PC?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry!

OpenStudy (welshfella):

you can use the calculator on that anyway the answer is 4.12 so you have 2 solutions -7 + 4.12 -7 - 4.12 -------- and --------- 2 2

OpenStudy (welshfella):

yw